Quantitative Problem: Jasper Jewelry has $150 million in sales. The company expects that its sales will increase 4% this year. Jasper's CFO uses a simple linear regression to forecast the company's inventory level for a given level of projected sales. On the basis of recent history, the estimated relationship between inventories and sales (in millions of dollars) is as follows:

Inventories = $11 + 0.09(Sales)

Given the estimated sales forecast and the estimated relationship between inventories and sales, what is your forecast of the company's year-end inventory level? Enter your answer in millions. For example, an answer of $10,550,000 should be entered as 10.55. Do not round intermediate calculations. Round your answer to two decimal places.
$   million
Answer #1

Inventory=11+0.09*Sales in millions of dollars=11+0.09*150*1.04=25.04 million

Year end inventory level is 25.04 million